# Equivariant Machine Learning on Graphs with Nonlinear Spectral Filters

### v1

@@ C.5 Diagonal Nonlinear Spectral Filter
Here, we present a new type of NLSFs, termed Diagonal NLSFs, which are index-by-index Index-NLSFs and band-by-band Value-NLSFs. We denote them as follows:
@@
where $a,e$ are as before. Note that these operators are also equivariant to our group actions.
We investigate Diagonal NLSFs in graph classification tasks as described in Sec. 5, including MUTAG, PTC, ENZYMES, PROTEINS, NCI1, IMDB-B, IMDB-M, and COLLAB datasets. Unlike Graph-NLSFs, which are fully spectral and map a sequence of frequency coefficients to an output vector, Diagonal NLSFs do not possess such a sequential spectral form. In Diagonal NLSFs, the index-by-index (or band-by-band) frequency response is projected back to the graph domain. Consequently, for graph classification tasks, we apply the readout function as defined for Pooling-NLSFs in Sec. 3.4.
Following App. C.3, we examine index-by-index Index and band-by-band Value settings using graph Laplacian $\mathbf{L}$ and normalized graph Laplacian $\mathbf{N}$, including $\Gamma_{\text{ind}}^{\mathtt{P}}(\mathbf{L},\cdot)$, $\Gamma_{\text{ind}}^{\mathtt{P}}(\mathbf{N},\cdot)$, $\Gamma_{\text{val}}^{\mathtt{P}}(\mathbf{L},\cdot)$, and $\Gamma_{\text{val}}^{\mathtt{P}}(\mathbf{N},\cdot)$, along with their variants using Laplacian attention, where $\mathtt{P}$ denotes the pooling function as in Tab. LABEL:tab:graph_classification_nlsfp_ablation_pooling.
Tab. LABEL:tab:graph_classification_nlsfp_diagonal_nlsf presents the graph classification accuracy using Diagonal NLSFs. It shows that incorporating Laplacian attention consistently improves classification performance, aligning with the findings in App. C.3. We note that Diagonal-NLSFs perform comparably to existing baselines in graph classification benchmarks. However, Diagonal-NLSFs are generally less effective compared to Pooling-NLSFs, as shown in Tab. LABEL:tab:graph_classification_nlsfp_ablation_pooling.
C.6 Uniform Sub-Bands

### v2

@@ F.6 Index-by-Index Index-NLSFs and Band-by-Band Value-NLSFs
Here, we present a new type of NLSFs: index-by-index Index-NLSFs and band-by-band Value-NLSFs. We denote them as follows:
@@
where $a,e$ are as before in Sec. 3.3. Note that these operators are also equivariant to our group actions.
We investigate index-by-index Index-NLSFs and band-by-band Value-NLSFs in graph classification tasks as described in Sec. 5, including MUTAG, PTC, ENZYMES, PROTEINS, NCI1, IMDB-B, IMDB-M, and COLLAB datasets. Unlike Graph-NLSFs, which are fully spectral and map a sequence of frequency coefficients to an output vector, index-by-index Index-NLSFs and band-by-band Value-NLSFs do not possess such a sequential spectral form. In index-by-index Index-NLSFs and band-by-band Value-NLSFs, the index-by-index (or band-by-band) frequency response is projected back to the graph domain. Consequently, for graph classification tasks, we apply the readout function as defined for Pooling-NLSFs in Sec. 3.4.
Following App. F.3, we examine index-by-index Index-NLSFs and band-by-band Value-NLSFs settings using graph Laplacian $\mathbf{L}$ and normalized graph Laplacian $\mathbf{N}$, including $\Gamma_{\text{ind}}^{\mathtt{P}}(\mathbf{L},\cdot)$, $\Gamma_{\text{ind}}^{\mathtt{P}}(\mathbf{N},\cdot)$, $\Gamma_{\text{val}}^{\mathtt{P}}(\mathbf{L},\cdot)$, and $\Gamma_{\text{val}}^{\mathtt{P}}(\mathbf{N},\cdot)$, along with their variants using Laplacian attention, where $\mathtt{P}$ denotes the pooling function as in Tab. LABEL:tab:graph_classification_nlsfp_ablation_pooling.
Tab. LABEL:tab:graph_classification_nlsfp_diagonal_nlsf presents the graph classification accuracy using index-by-index Index-NLSFs and band-by-band Value-NLSFs. It shows that incorporating Laplacian attention consistently improves classification performance, aligning with the findings in App. F.3. We note that index-by-index Index-NLSFs and band-by-band Value-NLSFs perform comparably to existing baselines in graph classification benchmarks. However, index-by-index Index-NLSFs and band-by-band Value-NLSFs are generally less effective compared to Pooling-NLSFs, as shown in Tab. LABEL:tab:graph_classification_nlsfp_ablation_pooling.
F.7 Node Classification Using Diag-NLSFs and Lead-NLSFs
Table 15: Semi-supervised node classification accuracy using NLSFs, diag-NLSFs, lead-NLSFs, and lead-diag-NLSFs.
Cora Citeseer Pubmed Chameleon Squirrel Actor
$\mathtt{att}$-Node-level diag-NLSFs 85.37$\pm$1.8 75.41$\pm$0.8 82.22$\pm$1.2 50.58$\pm$1.3 38.39$\pm$0.9 35.13$\pm$1.0
$\mathtt{att}$-Node-level NLSFs 86.03$\pm$1.2 74.87$\pm$1.2 83.15$\pm$1.5 50.12$\pm$1.2 36.23$\pm$1.6 35.01$\pm$1.7
$\mathtt{att}$-Node-level lead-NLSFs 85.26$\pm$0.4 74.16$\pm$1.4 82.24$\pm$0.9 49.86$\pm$1.9 34.21$\pm$1.1 33.68$\pm$1.1
$\mathtt{att}$-Node-level lead-diag-NLSFs 84.75$\pm$0.7 73.62$\pm$1.1 81.93$\pm$1.0 49.68$\pm$1.6 38.25$\pm$0.7 34.72$\pm$0.9
In App. C, we presented the diag-NLSFs, considering $\widetilde{d}=d$, and lead-NLSFs for leading filters that do not include orthogonal complements. We explore the leading filters, orthogonal complement, and diagonal operation in our NLSFs. The results of these investigations are summarized in Tab. LABEL:tab:all_nlsfs , which shows the node classification accuracy achieved using various configurations, including NLSFs, diag-NLSFs, lead-NLSFs, and lead-diag-NLSFs. We see that incorporating both the orthogonal complement and the multi-channel approach yields the highest classification accuracy.
F.8 Uniform Sub-Bands
@@
Appendix G Additional Related Equivariant GNNs
Equivariant GNNs are designed to handle graph data with symmetries. The output of an equivariant GNN respects the same transformation applied to the input. Depending on the application, the transformation could involve translations, reflections, or permutations, to name but a few [82, 46, 78]. Due to respecting the symmetries, equivariant GNNs can reduce model complexity and improve generalization [35, 27], which can be applied to test data and produce more interpretable representations [99, 51]. When symmetry plays a critical role, such as in physical simulations, molecular modeling, and protein folding, equivariant GNNs have been demonstrated to be effective [20]. For example, [91, 33] employ spherical harmonics to handle 3D molecular structures, and [82] simplify computations for explicit rotation and translation matrices by focusing on relative distances between nodes. In addition, [30] embeds symmetry transformations by parameterizing convolution operations over Lie groups, such as rotations, translations, and scalings, through Lie algebra.